OpenStudy (anonymous):

Find the surface area of this cone. Please help, I'm just stuck on one part of this question and I want to know what im doing wrong, thank you.

OpenStudy (anonymous):



OpenStudy (anonymous):

For the base I got =64 LA= I got 2376 SA=2440 but its wrong so idk what I did wrong

OpenStudy (anonymous):

A=πr(r+h2+r2)

OpenStudy (anonymous):

A=(3.14)(8)(8+19^2+8^2) A=(3.14)(8)(8+361+64) 3.13(8)(8+425) (3.14)(8)(433) (25.12)(433) ? Can you do it from here?

OpenStudy (anonymous):

that number seems extremely big to be the answer?

OpenStudy (anonymous):

(3.14)(8)(8+361+64) (3.14)(64+2888+512) (3.14(2888+576) 3.14(3400)

OpenStudy (anonymous):

oh wait when i copied the formula i didnt put in sqrt

OpenStudy (anonymous):

\[(3.14)(8)(8+\sqrt{361+64}\]

OpenStudy (anonymous):

\[(3.14)(8)(8+\sqrt{425}\]

OpenStudy (anonymous):

3016?

OpenStudy (anonymous):

(3.14))(8)(8+20.61) 3.14(64+164.9) 200.96+1319.2 1520.16

OpenStudy (anonymous):

that doesn't seem right

OpenStudy (anonymous):

(3.14)(8)(8+20.61) 25.12(8+20.61) 200.96+517.72 718.68

OpenStudy (anonymous):

is that the answer or do I plug that in and add that to the base?

OpenStudy (anonymous):

I used a calculator on google it said 719.19

OpenStudy (anonymous):

it was wrong :(

OpenStudy (anonymous):

719.19 or 718.68?

OpenStudy (anonymous):

the 719.19